OpenStudy (anonymous):

If r(x) = −4x − 1 and t(x) = 2x − 1, find (r + t)(7). −16 16 14 −14

OpenStudy (cggurumanjunath):

r(7)=-4(7)-1 t(7)=2(7)-1 (r+t)(7)=r(7)+t(7)

OpenStudy (cggurumanjunath):

now tell me what answer u got ?